1.) A company is producing a product that has the following data:

volume of sales per year = 3

00,000 units

selling price per unit = $10

0

variable cost = $6

0 per unit

fixed costs per year = $6

00,000

book depreciation = $4

00,000

tax depreciation = $50

0,000

debt interest = $1

00,000

tax rate = 40%

With the above data determine

(a) Before

-

tax profit

(b) After

-

tax profit

(c) Break

-

even yea

rly volume of sales

(d) Break

-

even selling price

(e) Plot the break

-

even chart
Answer #1

From the given data

Total sales revenue = selling price * total volume of shares

= 300,000*100

=30000,000 dollars

total cost of firm = variable cost + fixed cost

= variable price * total volume+ fixed cost

= 60* 300,000+600,000  

= 18600,000 dollars

Earning before interest tax depreciation(EBITD) = sales - cost

= 30000,000-18600,000

= 11400,000 dollars

Earnings before interest and tax (EBIT)= EBITD - depreciation

=11400,000-(400,000+500,000)

=10500,000 dollars

Earnings before tax = EBIT - interest

= 10500,000-100,00,

=10400,000 dollars

BEFORE TAX PROFIT = 10400,000 DOLLARS

tax = 40 % = 0.4* 10400,000 = 4160,000 dollars

AFTER TAX PROFIT = 10400,000-4160,000 = 6240,000 DOLLARS

Break even yearly volume of sales

= ( fixed cost )/ ( selling price per unit - variable cost per unit)

= 600000/(100-60)

=600000/40 = 15000 units

Break even selling price = break even quantity * selling price per unit

=15000*100=1500,000 dollars

PLOT

Ler the qpan-h be Sales Cioox) Break even--_ sales Break even Break Even chart Break Even vont